The core diameter of multimode step index fibre is 60 micrometer. The difference in
refractive index is 0.013. The core refractive index is 1.46. Determine the number of
guided modes when the operating wavelength is 0.75 micrometer.

Grade:12th pass

1 Answers

Manika gupta
3 years ago
Dear Student,
You may find your solution here.
=n1-n2/n1

n12_n22=(n1+n2)=(n1+n2/2)(n1_n2/n1)2n1

n1+n2/2n1

and

NA=n122

therfore

na=n12

v-no. or normalized frequence

v=n12_n22

No. of fibre traveling

for step index fibre

N=v2/2

for graded index fibre

N=v2/4

AskIITians expert.